मराठी
महाराष्ट्र राज्य शिक्षण मंडळएचएससी वाणिज्य (इंग्रजी माध्यम) इयत्ता १२ वी

In a town, 10 accidents take place in the span of 50 days. Assuming that the number of accidents follows Poisson distribution, find the probability

Advertisements
Advertisements

प्रश्न

In a town, 10 accidents take place in the span of 50 days. Assuming that the number of accidents follows Poisson distribution, find the probability that there will be 3 or more accidents on a day.

(Given that e-0.2 = 0.8187)

Solution:

Here, m = `square` and X − P(m) with parameter m.

The p.m.f. X is:

P(X = x) = `(e^(−m).m^x)/(x!), x = 0, 1, 2,...`

P(X ≥ 3) = 1 − P(X < 3)

= 1 − [`square + square + square`]

= `1 − [(e^(− 0.2)(0.2)^0)/(0!) + (e^(−0.2)(0.2)^1)/(1!) + (e^(−0.2)(0.2)^2)/(2!)]`

= 1 − [0.8187(1 + 0.2 + 0.02)]

= 1 − `square`

= `square`

रिकाम्या जागा भरा
बेरीज
Advertisements

उत्तर

Here, m = \[\boxed{0.2}\] and X − P(0.2) with parameter m.

The p.m.f. of X is X ∼ P(m).

P(X = x) = `(e^(−m)  .  m^x)/(x!), x = 0, 1, 2,...`

P(X ≥ 3) = 1 − P(X < 3)

= 1 − [\[\boxed{P(0)}\] + \[\boxed{P(1)}\] + \[\boxed{P(2)}\]]

= `1 − [(e^(− 0.2)(0.2)^0)/(0!) + (e^(−0.2)(0.2)^1)/(1!) + (e^(−0.2)(0.2)^2)/(2!)]`

= 1 − [0.8187(1 + 0.2 + 0.02)]

= 1 − \[\boxed{0.998814 }\]

= \[\boxed{0.001186}\]

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
2025-2026 (March) Board Question Paper

व्हिडिओ ट्यूटोरियलVIEW ALL [2]

संबंधित प्रश्‍न

The number of complaints which a bank manager receives per day is a Poisson random variable with parameter m = 4. Find the probability that the manager will receive -

(a) only two complaints on any given day.

(b) at most two complaints on any given day

[Use e-4 =0.0183]


 If a random variable X follows Poisson distribution such that P(X = l) =P(X = 2), then find P(X ≥ 1).  [Use e-2 = 0.1353] 


If X has a Poisson distribution with variance 2, find P (X = 4) 

[Use e-2 = 0.1353] 


If X has a Poisson distribution with variance 2, find P(X ≤ 4) 

[Use e-2 = 0.1353] 


If X has a Poisson distribution with variance 2, find 

Mean of X [Use e-2 = 0.1353] 


If X has Poisson distribution with m = 1, then find P(X ≤ 1) given e−1 = 0.3678


If X has Poisson distribution with parameter m and P(X = 2) = P(X = 3), then find P(X ≥ 2). Use e−3 = 0.0497


A car firm has 2 cars, which are hired out day by day. The number of cars hired on a day follows Poisson distribution with mean 1.5. Find the probability that (i) no car is used on a given day, (ii) some demand is refused on a given day, given e−1.5 = 0.2231.


It is known that, in a certain area of a large city, the average number of rats per bungalow is five. Assuming that the number of rats follows Poisson distribution, find the probability that a randomly selected bungalow has exactly 5 rats inclusive. Given e-5  = 0.0067.


It is known that, in a certain area of a large city, the average number of rats per bungalow is five. Assuming that the number of rats follows Poisson distribution, find the probability that a randomly selected bungalow has between 5 and 7 rats inclusive. Given e−5 = 0.0067.


If E(X) = m and Var(X) = m then X follows ______.


Solve the following problem :

If X follows Poisson distribution with parameter m such that
`("P"("X" = x + 1))/("P"("X" = x)) = (2)/(x + 1)`
Find mean and variance of X.


X : is number obtained on upper most face when a fair die is thrown then E(X) = ______


The expected value of the sum of two numbers obtained when two fair dice are rolled is ______.


Choose the correct alternative:

For the Poisson distribution ______


State whether the following statement is True or False:

X is the number obtained on upper most face when a die is thrown, then E(x) = 3.5


State whether the following statement is True or False:

If n is very large and p is very small then X follows Poisson distribution with n = mp


State whether the following statement is true or false:

lf X ∼ P(m) with P(X = 1) = P(X = 2) then m = 1.


If X – P(m) with P(X = 1) = P(X = 2), then find the mean and P(X = 2) given that e–2 = 0.1353

Solution: Since P(X = 1) = P(X = 2)

`(e^-mm^1)/(1!) = square`

∴ m = `square`

∴ mean = `square` = `square`

Then P(X = 2) = `square` = `square`


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×